[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jul 11, 2016

Dr. Craig A. Mattocks craig.mattocks at noaa.gov
Mon Jul 11 17:01:13 UTC 2016


Update on first HSU item: The Hurricane Specialists are reporting that the ATCF Atlantic desk workstation is now much more responsive after the upgrade to RHEL 6.8 and the deactivation of the aggressive “cachefilesd” (network data caching) daemon by NCO SysAdmin Brian Maher.

Craig

> On Jul 11, 2016, at 12:12 PM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:
> 
> Here is today's summary of NHC computer operations: 
> --- Monica
> 
> ----------
> 
> TAFB:
> TAFB submitted an NCO Helpdesk ticket for the printer in TAFB ops since it was failing to print anything including the TC Dvorak classifications. NCO has just fixed this as of 10 minutes ago.
> 
> HSU:
> -Brennan reported that a "cachefilesd" process was using all the CPU power over the weekend. This process was a consequence of the Red Hat upgrade and NCO has now disabled this process. 
> -The root cause of the Euro Ships guidance failing on Sat was due to their grib files arriving late and therefore were not decoded. This process worked again during the next run. 
> -Automated import of the TAFB Dvorak fixes into the ATCF have improved according to shift logs, but some additional tweaking is needed. TSB will continue to troubleshoot this.
> -A corrected SAB fix overwrote the original 12Z fix in the ATCF. TSB will attempt to grab the original and add it back into the fdeck. Additional troubleshooting may be required to permanently fix this if it becomes a repeat issue. 
> -HSU shift log emails are still occasionally sending duplicates. TSB will continue to monitor and troubleshoot this.
> -No FSSE trackers for the A shift on 7/10 at 6Z. TSB will investigate.
> -Forecasters reported that the Himawari data has not been available for the past few days. A series of issues with the dataflow from both NESDIS and NCO have affected the data reaching NHC. The dataflow has been restored, and the local processing code has caught up with the data stream so H8 is now available in NAWIPS again. See the SDM log and previous morning rounds for more info. 
> -The pac-atcf workstation is still slow. TSB and NCO will investigate.
> -The atlantic ATCF workstation was reported as being slow. It was reccomended that the R shift from 7/10 do a reboot to see if the performance improved. TSB has not heard anything further from HSU, so hopefully the reboot has improved the performance issues. 
> -HSU reported that the HSU ops printer has been printing black streaks and warning about the transfer cartridge being low. An NCO helpdesk ticket will need to be submitted for this. 
> -The hurricane hotline did not ring for the 7/10 17Z Medium Range call. TSB and NCO will investigate.
> 
> Other items:
> -Storm Surge AWIPS2 Sprint this week
> -NESDIS has been tasked to provide more transparency with the times available to request RSO operations. New changes to their website to indicate blackout or maintenance periods will be implemented in the coming months. See the SDM log for more info.
> 
> NHC RFCs:
> 1. NHC #182: Dave will implement new olor-coded dataflow visualization scripts for the NHC intranet today at 12Z
> 2. NCO #2098: GFDL v14.0.3 upgrade on 7/12 at 12Z, see TIN for more info
> 3. NCO #2101: HWRF v10.0.3 upgrade on 7/12 at 1330Z, see TIN for more info
> 
> 




More information about the Ncep.nhc.nco_contacts mailing list